PSEUDOROTATION IN TETRAHYDROFURAN: SUPPLEMENTARY DATA FROM THE FREE JET RO-TATIONAL SPECTRUM OF THE d_{8} SPECIES

\underline{WALTHER CAMINATI}, ASSIMO MARIS, SONIA MELANDRI, Dipartimento di Chimica ”G. Ciamician” dell’Università, Via Selmi 2, I-40126 Bologna, Italy; JUAN C. LÓPEZ, JOSÉ L. ALONSO, Departamento de Química-Física y Química-Inorgánica, Facultad de Ciencias, Universidad de Valladolid, E-47005 Valladolid, Spain; ROLF MEYER, Sonnenbergstrasse 18, Zufikon, CH-5621, Switzerland.

Tetrahydrofuran (THF) is a model molecule for the investigation of the hindered pseudorotation in saturated fivemembered rings. Two recent papers, based on millimeter-wave investigations, reported the vibrational spacings among the four lowest pseudorotation states of normal THF, and obtained contrastant results on the potential energy function: four equivalent minima^{a}, or 2 sets of two equivalent minima^{b}. In order to have more information on this problem, we planned to investigate the millimeter wave spectrum of the d_{8} species. We obtained a vibrational spacing of ca 15 GHz between the two lowest pseudorotation states: a result which is not in agreement with both the above mentioned potential energy functions.

aR. Meyer, J. C. López, J. L. Alonso, S. Melandri, W. Caminati, and P. G. Favero, J. Chem. Phys. 1999, 111, 7871.

bD. G. Melnik, S. Gopalakrishnan, T. A. Miller, F. C. D. Lucia, J. Chem. Phys. 2003, 118, 3589.